# CADSpotting: Robust Panoptic Symbol Spotting on Large-Scale CAD Drawings

### v2

@@ Abstract
We introduce CADSpotting, an efficient method for panoptic symbol spotting in large-scale architectural CAD drawings. Existing approaches struggle with the diversity of symbols, scale variations, and overlapping elements in CAD designs. CADSpotting overcomes these challenges by representing each primitive with densely sampled points, described by essential attributes like coordinates and colors. Building upon a unified 3D point cloud model for joint semantic, instance, and panoptic segmentation, CADSpotting learns robust feature representations. To enable accurate segmentation in large, complex drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining weighted voting and Non-Maximum Suppression (NMS). Moreover, we introduce a large-scale CAD dataset named LS-CAD to support our experiments. Each floorplan in LS-CAD has an average coverage of 1,000 m2 (versus 100m2 in the existing dataset), providing a valuable benchmark for symbol spotting research. Experimental results on FloorPlanCAD and LS-CAD datasets demonstrate that CADSpotting outperforms existing methods, showcasing its robustness and scalability for real-world CAD applications.
@@ 1 Introduction
Architectural Computer-Aided Design (CAD) drawings, especially those used for interior design, serve as detailed digital representations that convey essential information about a building’s structure, layout, and intricate details. These drawings include critical elements such as furniture placement, electrical layouts, and spatial organization, ensuring consistency and precision throughout the construction process. For instance, the Shanghai Mercedes-Benz Arena utilized over 3,000 CAD drawings to guide its structural design, optimize spatial arrangements, and streamline the construction workflow. As CAD designs continue to grow in complexity, automating the detection and interpretation of symbols within these floorplans becomes increasingly important in various downstream tasks, such as code compliance checking and 3D interior modeling [20].
@@
Traditional methods approached symbol recognition in a query-by-example manner [29], but these techniques struggle to handle the vast diversity of graphical symbols in real-world datasets. More recent learning-based approaches [25, 7] have shown significant progress in addressing the symbol spotting task. Early solutions [5, 27, 39] typically leverage convolutional neural networks (CNNs) or graph-based models to spot symbols. More advanced methods leverage Transformer architectures and attention mechanisms [6, 39], improving global relationship reasoning among symbols. These methods use rasterized pixel images derived from vector CAD drawings as input. Such conversion, however, introduces discrepancies between the original vector graphics and the rasterized images, resulting in errors due to the loss of precise geometric information.
@@
The representative work SymPoint [16] treats CAD drawings as sets of primary points, leveraging techniques from point cloud analysis for effective CAD symbol spotting. Such a representation simplifies model complexity while achieving strong performance on panoptic symbol spotting tasks in CAD drawings. Building on this foundation, SymPoint-V2 [17] introduces Layer feature-enhanced encoding to incorporate graphical layer information and a position-guided Training method to enhance model learning and accelerate convergence. Yet both approaches rely on manually defined primitive feature representations, i.e., four fixed primitive types. As a result, they struggle to capture the full diversity of CAD graphical primitives. In addition, these methods focus on processing small to medium-scale datasets due to the absence of efficient strategies for preserving connectivity during symbol spotting across expansive drawings. By far, there is still a lack of efficient methods to address the panoptic symbol spotting task in large-scale CAD drawings, as dense symbol clutter and significant variations in scale present substantial challenges for accurate segmentation and symbol recognition.
@@
In this paper, we present CADSpotting, a simple yet highly effective panoptic symbol spotting method tailored for handling CAD drawings at a much larger scale. Specifically, CADSpotting does not rely on fixed graphic primitive types. Instead, it densely samples points along CAD graphic primitives to construct comprehensive point cloud representations. Each point is represented by basic attributes, such as coordinates and color, forming a lightweight but expressive feature set. We utilize a unified 3D point cloud processing model to learn robust features from the sampled points. CADSpotting then employs a streamlined Transformer decoder module for panoptic symbol spotting, effectively leveraging the detailed learned features to achieve precise segmentation. To handle large-scale, real-world CAD drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining a weighted voting strategy with Non-Maximum Suppression (NMS) to enable accurate and efficient panoptic segmentation.
@@ 2 Related Work
Image segmentation is a fundamental task in computer vision, traditionally divided into two main categories: semantic segmentation [18, 2, 33, 1, 35] and instance segmentation [28, 8, 9, 13]. Semantic segmentation assigns a class label to every pixel in an image, while instance segmentation distinguishes individual objects within the same class. Recently, SAM [13] introduces a large model trained on over one billion masks, leveraging prompt-based interactions for enhanced interactive segmentation. SAM2 [24] extends this approach to video segmentation. However, semantic and instance segmentation methods struggle to accurately represent uncountable background elements, such as the sky or terrain. To address this challenge, Kirillov et al. [12] propose Panoptic Segmentation, which unifies semantic and instance segmentation, effectively handling both countable and uncountable background components. Early approaches [11, 31, 15, 3] primarily employ CNN-bases architectures, while recent advancements, such as Mask2Former [4], SegFormer [35], and OneFormer [9], leverage Transformer models, incorporating a mask attention mechanism to boost segmentation performance. Following the large-model paradigm of SAM, SEEM [40] further refines segmentation capabilities using a prompt-driven design. Despite these advances, most existing segmentation models remain pixel-centric, limiting their effectiveness in processing vector graphics like CAD drawings. Consequently, they often fail to capture critical vector-based information, such as overlapping relationships and precise geometric attributes.
@@
Early efforts [22, 23, 26, 10, 36] in symbol spotting on CAD drawings rely on handcrafted feature descriptors, such as shape and structure, coupled with techniques like sliding window searches or graph matching for symbol retrieval. The introduction of deep learning techniques marked a significant shift, with models based on Faster R-CNN [25] and YOLO v3 [7] greatly enhancing the accuracy of symbol recognition tasks [27]. However, traditional symbol spotting methods are primarily designed for countable objects and do not effectively handle uncountable symbols, which are common in CAD environments. To address this limitation, Fan et al. [5] introduce the FloorPlanCAD dataset and CNN-GCN method, enabling the recognition of both countable and uncountable symbols for more comprehensive CAD parsing. Building on this, Zheng et al. [39] propose GAT-CADNet, a model that represents CAD drawings as graph structures and performs segmentation by predicting adjacency matrices. Further advancements include CADTransformer [6], which utilizes the Vision Transformer to extract features from scalar maps, leading to improved predictions. Meanwhile, significant progress has been made in the field of 3D point cloud segmentation [38, 34, 14, 30]. The seminal work SymPoint [16] proposes a novel method that converts symbols into primary point representations, thereby enhancing feature extraction through point cloud analysis and upsampling techniques. Concurrently, Liu et al. [17] extend SymPoint by incorporating LFE and PGT modules, further improving feature representation and model performance. Despite these advances, the manually defined feature representation is often constrained by a limited set of graphical primitive types, which hinders their ability to capture the full diversity of CAD symbols.
@@ 3 Method
An overview of our CADSpotting method is provided in Fig.2. By integrating a dense point sampling strategy within a unified point cloud processing model, our approach efficiently extracts features for CAD primitives, producing a robust representation that supports various primitive types. Details are provided in Sec. 3.1. Then, a straightforward transformer decoder is applied to perform semantic, instance, and panoptic segmentation, with the corresponding loss function detailed in Sec.3.2. Additionally, we propose a novel Sliding Window Aggregation technique in Sec.3.3. Using a weighted voting strategy and Non-Maximum Suppression (NMS), this method allows our method to process expansive CAD drawings and perform large-scale panoptic segmentation.
@@ 3.1 Primitive Feature Learning
SymPoint [16] introduces a new concept that represents CAD primitives as a set of 2D points. It constructs a handcrafted feature for each graphical primitive, capturing details such as type (e.g., line, arc, circle, ellipse), length, etc. SymPoint then applies point cloud analysis to learn primitive representations, further enhancing connectivity through an Attention with Connection Module. A masked-attention transformer decoder is used to complete the panoptic symbol spotting task. Although SymPoint demonstrates strong performance in CAD symbol spotting, its feature representation is manually defined and limited to four fixed primitive types, which may constrain its capacity to fully capture the diverse range of CAD symbols. To address these constraints, we propose a novel graphical primitive feature representation for CAD drawings that incorporates dense point sampling within a unified point cloud analysis model to improve both the efficiency and generalizability of primitive-level representations.
@@
In CAD drawings, graphical primitives are typically categorized into line types and shape types. Line primitives encompass straight segments, curves, and complex contours such as Bézier curves, while shape primitives include basic geometric forms like rectangles, circles, and ellipses. Inspired by SymPoint [16], we propose a dense point sampling strategy that performs equidistant sampling on each primitive to generate dense point data for feature extraction. Rather than assigning fixed types to CAD graphical primitives and using handcrafted features to represent them, our dense point sampling method offers greater efficiency and adaptability, making it suitable for a wide range of CAD primitives. Fig.3 illustrates the process of our dense point sampling method.
Given a CAD drawing containing primitives, we perform dense sampling to generate a 2D point cloud from these primitives at a fixed distance . Each 2D point is represented by a six-dimensional vector :
where represent the 2D coordinates of the point, is set to 0, and represent the color information (red, green, blue channels) of the point. We then employ Point Transformer V3 [34] (PTv3) as the backbone for feature extraction, leveraging its design optimized for the unordered nature of point clouds, which provides high efficiency and robust performance in feature extraction.
@@
To obtain the final feature representation for CAD graphical primitives, we apply primitive mixed pooling to transform point-wise features into primitive-wise features. Specifically, we perform pooling on densely sampled points within each primitive , aggregating the point features of each primitive into a single primitive-wise feature vector:
where denotes the total number of 2D sampled points, and represents the number of primitives. Let represent a single primitive with its corresponding dense point cloud . The -th dimension feature of is determined by the sum of the maximum and average values of the corresponding -th dimension features across all points :
This pooling method condenses a large set of dense point features into a more compact set of primitive features, effectively reducing computational complexity and storage requirements. At the same time, primitive-wise pooling retains the essential information of the contained points, preserving global consistency in the primitive representation and minimizing information loss. Loss calculations are also conducted at the primitive level, aligning with our final goal of performing CAD spotting tasks.
@@ 3.2 Panoptic Symbol Spotting
Following OneFormer3D [14], we partition the decoder output into two parts: the first outputs represent instance symbol spotting proposals, and the subsequent outputs correspond to semantic symbol spotting. This configuration enables panoptic symbol spotting by integrating instance and semantic information for comprehensive scene understanding. Here, is manually defined, and corresponds to the number of semantic categories.
We translate decoder outputs to semantic proposals and instance proposals. A semantic proposal includes the log-likelihood values for each primitive being classified into different semantic labels and can be further processed through thresholding to generate a binary mask. Similarly, an instance proposal also contains log-likelihood values, but these indicate the likelihood of each primitive being classified as a specific instance. During inference, we select the top- instance proposals of the highest confidence scores and apply matrix-NMS [32] to suppress non-maximal instance predictions.
Loss Function.
The loss function is composed of classification loss and primitive mask loss. The classification loss is defined as a multi-class cross-entropy loss. The primitive mask loss combines binary cross-entropy loss, , with the Dice loss [21], , to evaluate the correspondence between each predicted instance and its ground truth counterpart.
@@ 3.3 Sliding Window Aggregation
To address the challenge of panoptic symbol segmentation at large scales, a key issue is ensuring the generalization of instance symbol segmentation across varying scales. During inference on larger drawings than those encountered during training, the network must produce significantly more instance proposals. However, this process is limited by the fixed number of instance queries inherent in Mask Transformer-based methods, a constraint that also affects our approach. Nonetheless, CAD instance symbol segmentation offers certain advantages over general image instance segmentation. Specifically, in CAD floorplans, instances of the same category maintain relatively consistent sizes across different scenes when a uniform scale is applied. Consequently, we can reasonably assume that the size of instances in inference will not exceed that of the same category in the training images.
@@
For instance proposals, we directly apply Non-Maximum Suppression (NMS) to the instance proposals obtained from all windows. A potential instance may be partially or fully observed across multiple windows. By ensuring a sufficiently small sliding step size, each instance is fully captured within at least one window. Incomplete proposals from other windows are subsequently filtered out during the NMS process.

### v3

@@ Abstract
We introduce CADSpotting, an effective method for panoptic symbol spotting in large-scale architectural CAD drawings. Existing approaches struggle with symbol diversity, scale variations, and overlapping elements in CAD designs. CADSpotting overcomes these challenges by representing primitives through densely sampled points with attributes like coordinates and colors, using a unified 3D point cloud model for robust feature learning. To enable accurate segmentation in large, complex drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining weighted voting and Non-Maximum Suppression (NMS). Moreover, we introduce LS-CAD, a new large-scale CAD dataset to support our experiments, with each floorplan covering around 1,000 m2, significantly larger than previous benchmarks. Experiments on FloorPlanCAD and LS-CAD datasets show that CADSpotting significantly outperforms existing methods. We also demonstrate its practical value through automating parametric 3D reconstruction, enabling interior modeling directly from raw CAD inputs.
@@ 1 Introduction
Architectural Computer-Aided Design (CAD) drawings, especially those used for interior design, serve as detailed digital representations that convey essential information about a building’s structure, layout, and intricate details. These drawings include critical elements such as furniture placement, electrical layouts, and spatial organization, ensuring consistency and precision throughout the construction process. For instance, the Shanghai Mercedes-Benz Arena utilize over 3,000 CAD drawings to guide its structural design, optimize spatial arrangements, and streamline the construction workflow. As CAD designs continue to grow in complexity, automating the detection and interpolation of symbols within these floorplans becomes increasingly important in various downstream tasks, such as code compliance checking and 3D interior modeling [20].
@@
Traditional methods approach symbol recognition in a query-by-example manner [29], but these techniques struggle to handle the vast diversity of graphical symbols in real-world datasets. More recent learning-based approaches [25, 7] have shown significant progress in addressing the symbol spotting task. Early solutions [5, 27, 40] typically leverage convolutional neural networks (CNNs) or graph-based models to spot symbols. More advanced methods leverage Transformer architectures and attention mechanisms [6, 40], improving global relationship reasoning among symbols. These methods use rasterized pixel images derived from vector CAD drawings as input. Such conversion, however, introduces discrepancies between the original vector graphics and the rasterized images, resulting in errors due to the loss of precise geometric information.
@@
The representative work SymPoint [16] treats CAD drawings as sets of primary points, leveraging techniques from point cloud analysis for effective CAD symbol spotting. Such a representation simplifies model complexity while achieving strong performance on panoptic symbol spotting tasks in CAD drawings. Building on this foundation, SymPoint-V2 [17] introduces Layer feature-enhanced encoding to incorporate graphical layer information and a position-guided Training method to enhance model learning and accelerate convergence. Yet both approaches rely on manually defined primitive feature representations, i.e., four fixed primitive types. As a result, they struggle to capture the full diversity of CAD graphical primitives. In addition, these methods focus on processing small to medium-scale datasets due to the absence of effective strategies for preserving connectivity during symbol spotting across expansive drawings. By far, there is still a lack of feasible methods to address the panoptic symbol spotting task in large-scale CAD drawings, as dense symbol clutter and significant variations in scale present substantial challenges for accurate segmentation and symbol recognition.
@@
In this paper, we present CADSpotting, a simple yet highly effective panoptic symbol spotting method tailored for handling CAD drawings at a much larger scale. Specifically, CADSpotting does not rely on fixed graphic primitive types. Instead, it densely samples points along CAD graphic primitives to construct comprehensive point cloud representations. Each point is represented by basic attributes, such as coordinates and color, forming a lightweight but expressive feature set. We utilize a unified 3D point cloud processing model to learn robust features from the sampled points. CADSpotting then employs a streamlined Transformer decoder for panoptic symbol spotting, effectively leveraging the learned features to achieve precise segmentation. To handle large-scale, real-world CAD drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining a weighted voting strategy with Non-Maximum Suppression (NMS) to enable accurate and effective panoptic segmentation.
@@
We also demonstrate the effectiveness of CADSpotting by automating the generation of 3D interior models through parametric modeling. Using precise instance-level data obtained from CADSpotting, we compute spatial parameters for architectural elements, including wall contours, door and window positions, door orientations, and pivot points. With these spatial parameters clearly defined, we leverage parametric reconstruction techniques within Blender to efficiently build 3D interior models directly from raw CAD data. This approach significantly streamlines the modeling process, enhancing productivity and accuracy. To the best of our knowledge, our proposed pipeline is the first comprehensive solution that seamlessly integrates CAD segmentation with automated 3D interior model reconstruction.
@@ 2 Related Work
Image segmentation is a fundamental task in computer vision, traditionally divided into two main categories: semantic segmentation [18, 2, 34, 1, 36] and instance segmentation [28, 8, 9, 13]. Semantic segmentation assigns a class label to every pixel in an image, while instance segmentation distinguishes individual objects within the same class. Recently, SAM [13] introduces a large model trained on over one billion masks, leveraging prompt-based interactions for enhanced interactive segmentation. SAM2 [24] extends this approach to video segmentation. However, semantic and instance segmentation methods struggle to accurately represent uncountable background elements, such as the sky or terrain. To address this challenge, Kirillov et al. [12] propose Panoptic Segmentation, which unifies semantic and instance segmentation, effectively handling both countable and uncountable background components. Early approaches [11, 31, 15, 3] primarily employ CNN-bases architectures, while recent advancements, such as Mask2Former [4], SegFormer [36], and OneFormer [9], leverage Transformer models, incorporating a mask attention mechanism to boost segmentation performance. Following the large-model paradigm of SAM, SEEM [41] further refines segmentation using a prompt-driven design. Despite these advances, most existing segmentation models remain pixel-centric, limiting their effectiveness in processing vector graphics like CAD drawings. Consequently, they often fail to capture critical vector-based information, such as overlapping relationships and precise geometric attributes.
@@
Early efforts [22, 23, 26, 10, 37] in symbol spotting on CAD drawings rely on handcrafted feature descriptors, such as shape and structure, coupled with techniques like sliding window searches or graph matching for symbol retrieval. The introduction of deep learning techniques marks a significant shift, with models based on Faster R-CNN [25] and YOLO v3 [7] greatly enhancing the accuracy of symbol recognition tasks [27]. However, traditional symbol spotting methods are primarily designed for countable objects and do not effectively handle uncountable symbols, which are common in CAD environments. To address this limitation, Fan et al. [5] introduce the FloorPlanCAD dataset and CNN-GCN method, enabling the recognition of both countable and uncountable symbols for more comprehensive CAD parsing. Countable object instances (e.g., windows, doors) can be clearly identified and individually counted whereas uncountable stuff (e.g., wall, railing) refers to elements that cannot be individually distinguished. Building on this, Zheng et al. [40] propose GAT-CADNet, a model that represents CAD drawings as graph structures and performs segmentation by predicting adjacency matrices. Further advancements include CADTransformer [6], which utilizes the Vision Transformer to extract features from scalar maps, leading to improved predictions. Meanwhile, significant progress has been made in the field of 3D point cloud segmentation [39, 35, 14, 30]. The seminal work SymPoint [16] proposes a novel method that converts symbols into primary point representations, thereby enhancing feature extraction through point cloud analysis and upsampling techniques. Concurrently, Liu et al. [17] extend SymPoint by incorporating LFE and PGT modules, further improving feature representation and model performance. Despite these advances, the manually defined feature representation is often constrained by a limited set of graphical primitive types, which hinders their ability to capture the full diversity of CAD symbols.
@@ 3 Method
An overview of our CADSpotting method is illustrated in Fig.2. By incorporating a dense point sampling strategy within a unified point cloud processing model, our method effectively extracts features for CAD primitives, yielding a robust representation that supports various primitive types. We describe each module in detail in the following sections.
@@ 3.1 Primitive Feature Learning
SymPoint [16] introduces a new concept that represents CAD primitives as a set of 2D points. It constructs a handcrafted feature for each graphical primitive, capturing details such as type (e.g., line, arc, circle, ellipse), length, etc. SymPoint then applies point cloud analysis to learn primitive representations, further enhancing connectivity through an Attention with Connection Module. A masked-attention transformer decoder is used to complete the panoptic symbol spotting task. Although SymPoint shows strong performance in CAD symbol spotting, its feature representation is manually defined and limited to four fixed primitive types, which may constrain its capacity to capture the diverse range of CAD symbols. Real-world CAD drawings often contain various complex elements, such as Bézier curves, rather than simple predefined line types. Therefore, methods that rely on handcrafted features lack sufficient robustness. To address these constraints, we propose a novel graphical primitive feature representation for CAD drawings that incorporates dense point sampling within a unified point cloud analysis model to improve generalizability of primitive-wise representations.
@@
In CAD drawings, graphical primitives are typically categorized into line types and shape types. Line primitives encompass straight segments, curves, and complex contours such as Bézier curves, while shape primitives include basic geometric forms like rectangles, circles, and ellipses. Inspired by SymPoint [16], we propose a dense point sampling strategy that performs equidistant sampling on each primitive to generate dense point data for feature extraction. Rather than assigning fixed types to CAD graphical primitives and using handcrafted features to represent them, our dense point sampling method offers greater feasibility and adaptability, making it suitable for a wide range of CAD primitives. Fig.3 illustrates the process of our dense point sampling method.
Given a CAD drawing containing primitives, we perform dense sampling to generate a 2D point cloud from these primitives at a fixed distance . Each 2D point is represented by a six-dimensional vector: ,where represent the 2D coordinates of the point, is set to 0, and represent the color information (red, green, blue channels) of the point. We then employ Point Transformer V3 [35] (PTv3) as the backbone for feature extraction, leveraging its design optimized for the unordered nature of point clouds, which provides robust performance in feature extraction.
@@
To obtain the final feature representation for CAD primitives, we apply Primitive Mixed Pooling (PMP) to transform point-wise features into primitive-wise features. Let denote our pooling operator, which aggregates the point features into primitive-wise feature vector :
where denotes the total number of sampled points, and represents the number of primitives. Let represent the single primitive with its corresponding dense point cloud . The -th dimension feature of is determined by the sum of the maximum and average values of the corresponding -th dimension features across all points:
This pooling method condenses a large set of dense point features into a more compact set of primitive-level features, effectively reducing computational complexity and storage requirements. At the same time, primitive-level pooling retains the essential information of the contained points, preserving global consistency in the primitive representation and minimizing information loss. Loss calculations are also conducted at the primitive-level, aligning with our final goal of performing CAD spotting tasks.
@@ 3.2 Panoptic Symbol Spotting
Following OneFormer3D [14], we partition the decoder output into two parts: the first represent instance proposals, and the subsequent outputs correspond to semantic symbol spotting. The output dimension of the + vector is , integrating both instance and semantic information. The first dimensions are used to generate instance masks, while the remaining dimensions determine the semantic category of each primitive. This configuration enables panoptic symbol spotting by integrating instance and semantic information for comprehensive scene understanding.
We translate decoder outputs to semantic proposals and instance proposals. A semantic proposal includes the log-likelihood values for each primitive being classified into different semantic labels and can be further processed through thresholding to generate a binary mask. Similarly, an instance proposal also contains log-likelihood values, but these indicate the likelihood of each primitive being classified as a specific instance. Additionally, we employ a bipartite matching strategy based on the Hungarian algorithm to find the optimal correspondence between predicted results and ground truth labels while training. During inference, we select the top- instance proposals of the highest confidence scores and apply matrix-NMS [33] to suppress non-maximal instance predictions.
Loss Function. The loss function is composed of classification loss and primitive mask loss. The classification loss is defined as a multi-class cross-entropy loss. The primitive mask loss combines binary cross-entropy loss, , with the Dice loss [21], , to evaluate the correspondence between predicted instances and ground truth.
@@ 3.3 Sliding Window Aggregation
To address the challenge of panoptic symbol segmentation at large scales, a key issue is ensuring the generalization of instance symbol segmentation across varying scales. During inference on larger drawings than those encountered during training, the network must produce significantly more instance proposals. However, this process is limited by the fixed number of instance queries inherent in Mask Transformer-based methods, a constraint that also affects our approach. Nonetheless, CAD instance segmentation offers certain advantages over general image instance segmentation. Specifically, in CAD floorplans, instances of the same category maintain relatively consistent sizes across different scenes when a uniform scale is applied. Consequently, we can reasonably assume that the size of instances in inference will not exceed that of the same category in the training images.
@@
For instance proposals, we utilize Sparse Non-Maximum Suppression(Sparse-NMS) on the proposals generated from all windows, which increases processing efficiency (detailed analysis is provided in the supplementary material). A potential instance may be partially or fully observed across multiple windows. By ensuring a sufficiently small sliding step size, each instance is fully captured within at least one window. Incomplete proposals from other windows are subsequently filtered out during the NMS process.
3.4 Automated 3D Interior Reconstruction
CADSpotting delivers precise panoptic segmentation results, serving as the basis for our parametric 3D interior modeling algorithm. Utilizing detailed instance information and primitive positions from segmentation, we apply instance-level processing techniques to compute spatial parameters for key architectural elements, specifically walls, doors, and windows. For doors, we determine their positions, orientations, and pivot points by analyzing the relationships between arcs and lines within each segmented instance. Window positions, on the other hand, are directly extracted from their corresponding instances. Complex wall structures require a specialized approach: we first traverse and merge endpoints of adjacent lines within wall instances, forming closed polygons. These polygons are then rasterized into binary images, from which final wall contours are identified using the method from [32]. With these spatial parameters, we reconstruct a complete 3D interior model using a parametric reconstruction pipeline in Blender. The necessary 3D primitives for walls, doors, and windows are manually prepared in Blender, providing reusable templates for streamlined parametric interior modeling. This approach significantly enhances efficiency, automating the conversion of raw CAD data into structurally accurate 3D interior models within minutes through procedural generation. An example result from our automated 3D interior reconstruction method is illustrated in Fig. 4, with additional details provided in the supplementary material.